OpenStudy (anonymous):

What is the slope of the line passing through the point (3, −4) and perpendicular to the line y = 3x + 5? Write your answer as a fraction.

OpenStudy (anonymous):

Is it -1/3 ?

OpenStudy (anonymous):

perpendicular means the negative inverse. 3 is the slope of the original. So yes, -1/3 is the slope of the line perpendicular to it.

OpenStudy (anonymous):

perpendicular means product of theire slopes = -1 ie m1 *m2 = -1

OpenStudy (anonymous):

y = 3x + 5 Slope = 3 3 * m2 = -1 m2 = -1/3